a. slow in the internal 6. Show that the integral li converges.
Answer #1

& salin = we hace to show snel du I Jurph is converges. e t c) (cena prom) = 1 - 2 + c } = -20 + 4 + 2.83 x 3 +2.75 0+225-2 sSince I dn y convergent is also comersant Hance thence of Treen a converses ů convergess